# Beyond Two-Stage Training: Cooperative SFT and RL for LLM Reasoning

### v1

@@ Abstract
Reinforcement learning (RL) has proven effective in incentivizing the reasoning abilities of large language models (LLMs), but suffers from severe efficiency challenges due to its trial-and-error nature. While the common practice employs supervised fine-tuning (SFT) as a warm-up stage for RL, this decoupled two-stage approach limits interaction between SFT and RL, thereby constraining overall effectiveness. This study introduces BRIDGE, a novel method for learning reasoning models that employs bilevel optimization to facilitate better cooperation between these training paradigms. By conditioning the SFT objective on the optimal RL policy, our approach enables SFT to meta-learn how to guide RL’s optimization process. During training, the lower-level performs RL updates while simultaneously receiving SFT supervision, while the upper-level explicitly maximizes the cooperative gain—the performance advantage of joint SFT-RL training over RL alone. Empirical evaluations across three LLMs and five reasoning benchmarks demonstrate that our method consistently outperforms baselines and achieves a better balance between effectiveness and efficiency. [^3] The code will be available at https://github.com/ChanLiang/BRIDGE
@@ 1 Introduction
The emergence of OpenAI’s o1 (OpenAI, ) and DeepSeek-R1 (DeepSeek-AI et al., 2025) marks a significant advance in LLM reasoning capabilities, particularly for challenging tasks such as mathematics (Cobbe et al., 2021; Hendrycks et al., 2021b) and programming (Chen et al., 2021; Codeforces, 2025). The key technique driving this progress is large-scale, rule-based RL. However, the inherently trial-and-error nature of RL renders the training process highly inefficient. An alternative approach is SFT on curated long chain-of-thought (CoT) datasets, which enables models to rapidly acquire effective reasoning patterns through imitation learning. While more sample-efficient, SFT typically exhibits poorer performance and generalization than RL (Chu et al., 2025).
In practice, production-scale training pipelines often adopt a multi-stage paradigm, using SFT as a warm-up phase before applying RL. For example, DeepSeek-R1 (DeepSeek-AI et al., 2025) undergoes multiple rounds of SFT and RL to refine reasoning performance. However, our experiments reveal that this approach fails to fully leverage the strengths of both methods. The fully decoupled two-stage setup can suffer from catastrophic forgetting and inefficient exploration. This raises a natural question:
@@
To explore this possibility, we first propose a simple baseline that alternates between SFT and RL updates during training. Despite its simplicity, this approach improves both convergence efficiency and final performance. However, such independent updates cannot guarantee improvements over RL alone, as not all SFT updates benefit RL optimization. Building on this insight, we develop BRIDGE, a cooperative learning framework based on bilevel optimization, where SFT serves as the upper-level problem and RL as the lower-level problem. By solving this nested structure—with the SFT objective explicitly conditioned on the RL solution—SFT provides targeted guidance that directly supports RL’s optimization process.
@@
To validate the effectiveness of our approach, we conduct experiments with three LLMs across five diverse benchmark datasets covering both standard and competition-level math reasoning tasks. Results demonstrate that BRIDGE consistently outperforms all baselines—including SFT, RL-zero, cold-start, and our naive alternating baseline—while requiring less wall-clock training time. These improvements confirm the benefits of tightly coupling SFT and RL through bilevel optimization rather than treating them as separate phases.
@@ 2.1 Fine-tuning Methods for Reasoning Models
We consider a large language model (LLM) parameterized by $\bm{\theta}$, which defines a conditional distribution $\pi(y|x;\bm{\theta})$ over output sequences $y$ given input sequences $x$. This work focuses on three widely used methodologies for fine-tuning $\bm{\theta}$ to enhance the model’s reasoning capabilities.
@@
Rule-based Reinforcement Learning. Reinforcement learning with verifiable rewards has gained increasing attention for its effectiveness in training advanced reasoning models such as DeepSeek-R1 (DeepSeek-AI et al., 2025). Given a dataset $\mathcal{D}_{\mathrm{RL}}:=\{(x,y)\}$ with verifiable outputs—such as mathematics competition problems—the objective of rule-based RL is formulated as:
@@
Here, $y$ denotes the ground-truth answer and $\hat{y}$ is the model’s predicted output. The equivalence relation $\hat{y}\equiv y$ is typically computed by a domain-specific verifier (e.g., a symbolic math engine). This objective is commonly solved using policy optimization methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (DeepSeek-AI et al., 2025).
@@ 2.3 A Simple Alternating Baseline
As shown in Figure 1, this alternating strategy converges faster than pure RL and achieves better final performance than both standalone SFT and two-stage cold-start training. While this integration yields empirical gains, the current formulation treats SFT and RL as independent update processes with no guarantee that alternating updates will consistently outperform RL method alone. This limitation raises a natural question: How can we design training strategies that ensure better cooperation between SFT and RL leads to guaranteed superior performance compared to standalone RL?
@@ 3.1 BRIDGE: Cooperative Meta-Learning for SFT and RL
Given an SFT dataset $\mathcal{D}_{\mathrm{SFT}}$ and an RL dataset $\mathcal{D}_{\mathrm{RL}}$ (defined in Section 2.1), our objective is to integrate policy optimization (Eq. (2)) with supervised learning (Eq. (1)). We propose the following cooperative meta-learning formulation:
@@
For clarity, we express Equation (4) in simplified notation:
@@
The augmented model architecture, comprising base model parameters $\theta$ and LoRA parameters $w$, is essential for enabling cooperative learning. This separation allows the upper- and lower-level objectives to co-adapt during training, as illustrated in Figure 2. Without this architectural separation, our formulation (Equation (4)) would collapse to a Model-Agnostic Meta-Learning (MAML)-style setup (Finn et al., 2017), where the lower-level solution reduces to a single gradient step used to update the upper-level SFT parameters. In this case, RL learning is disabled, and the cooperation between SFT and RL is lost.
@@ 3.2 Learning Algorithm
To solve the bilevel optimization problem in Eq. (5), we employ penalty-based methods (Shen & Chen, 2023; Shen et al., 2025) to avoid expensive second-order derivative computations. We first reformulate (5) as a single-level problem amenable to efficient first-order optimization.
@@
The overall algorithm of BRIDGE is presented in Algorithm 2.
@@ 3.3 Intuition Behind the Update Rules
Lower-level update: Curriculum-weighted gradient fusion. The update rule for $\theta$ in Eq. (8) performs a convex combination of SFT and RL gradients. As $\lambda$ increases from 0 to 1 during training, the algorithm smoothly transitions from pure imitation learning to pure reinforcement learning.
This adaptive curriculum (Bengio et al., 2009) reflects the model’s evolving capabilities: early in training, when the base model lacks strong reasoning abilities, it benefits primarily from imitating expert demonstrations. As the model develops competence in generating correct solutions, it can increasingly leverage reward signals through exploration, making RL updates progressively more valuable.
Upper-level update: Maximizing cooperative gain. The update for $w$ in Eq. (11) solves the bilevel problem by finding LoRA parameters $w$ that ensure the RL-optimized model $\theta^{*}(w)$ also excels on the supervised dataset $\mathcal{D}_{\mathrm{SFT}}$.
The update in Eq. (11) can be interpreted as performing gradient ascent on the following objective:
@@ 4.1 Settings
We use two datasets for RL training: LIMR (Li et al., 2025) containing 1.3k unique problems and MATH (Hendrycks et al., 2021a) with 8.5k problems. For the SFT dataset, we pair queries from LIMR and MATH with corresponding intermediate reasoning traces extracted from DeepSeekMath-103k (He et al., 2025), which were distilled from the DeepSeek-R1 model. We evaluate on five mathematical reasoning benchmarks: MATH500 (Hendrycks et al., 2021a), Minerva Math (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), AIME 2024, and AMC 2023.
@@
To demonstrate the generality of our approach, we experiment with three LLMs: Qwen2.5-3B (Yang et al., 2024), Llama-3.2-3B-Instruct (Grattafiori et al., 2024), and Qwen2-8B-Base (Yang et al., 2025). All models use prompt formats consistent with SimpleRL (Zeng et al., 2025).
Reward Function.
Following SimpleRL (Zeng et al., 2025), we employ a binary reward based on answer correctness: +1 for correct final answers and 0 otherwise. We deliberately exclude format-based rewards, which can constrain exploration and reduce performance, particularly for base models.
Implementation Details.
[…]
@@ 4.2 Baselines
We compare BRIDGE against five baselines on the same base architectures:
@@
The base model or its instruction-tuned variant without additional reasoning-specific training, serving as performance lower bounds.
@@
Models trained exclusively on curated reasoning traces without reinforcement learning, demonstrating the capabilities and limitations of pure imitation learning.
@@
Reinforcement learning applied directly to the base model without prior fine-tuning, evaluating the effectiveness of exploration from scratch.
@@
A two-stage pipeline with SFT pretraining followed by RL fine-tuning, where phases are fully decoupled with no interaction between objectives.
@@
We introduce this baseline as an ablation study, which alternates between SFT and RL updates without the cooperative optimization. Despite being simple and effective, this straightforward independent alternating optimization approach allows us to isolate the additional gains from BRIDGE’s cooperative mechanism.
@@ 4.3 Experimental Results
Naive Alter. 65.2 25.3 27.1 6.7 42.5 33.4 (+3.7)
BRIDGE 66.2 23.9 28.9 13.3 47.5 36.0 (+11.8)
Table 1: Performance of BRIDGE compared to baseline methods across five math benchmarks. Average performance improvements (%) over Cold-start are highlighted in blue.
Generalization to benchmarks. We evaluate the generalization ability of BRIDGE across five diverse mathematical reasoning benchmarks. As shown in Table 1, BRIDGE consistently outperforms baseline methods, achieving consistently accuracy improvements on Minerva Math, Olympiad Bench, AIME24, and AMC23. Overall, BRIDGE yields an average improvement of 11.8% over RL-zero and Cold-start, highlighting its effectiveness and robustness across tasks of varying difficulty.
Baseline methods tend to yield larger improvements on relatively easier benchmarks but generalize poorly to more complex reasoning tasks. For example, the Cold-start method underperforms RL-zero on Minerva Math, Olympiad Bench, and AMC23, potentially due to overfitting during the prior SFT phase. While the Naive Alternative partially mitigates this issue—maintaining performance on harder benchmarks—its gains remain limited. In contrast, BRIDGE achieves consistent and substantial improvements on the more challenging benchmarks. These results underscore BRIDGE’s superior generalizability in handling complex mathematical reasoning.
@@
Table 2: Performance on Llama3.2-3B-Instruct.
@@
Table 3: Performance on Qwen3-8B-Base.
Effectiveness across LLMs. We expand our experiments to additional LLMs: Qwen3-8B-Base and Llama3.2-3B-Instruct. As shown in Tables 3 and 2, BRIDGE consistently outperforms all baselines across diverse architectures. On Qwen3-8B-Base, BRIDGE achieves 16.3% improvement over RL-zero and 9.7% over Cold-start. On Llama3.2-3B-Instruct, gains are more pronounced with 13.5% over RL-zero and 30.9% over Cold-start. These results demonstrate BRIDGE’s robust effectiveness across different model families and training configurations.
Performance on varied fine-tuning epochs. We assess BRIDGE’s effectiveness across different fine-tuning epochs on Qwen2.5-3B using average performance across epochs as the metric. As shown in Table 4, BRIDGE achieves the highest average performance.
Among the baselines, Cold-start yields the second-best trade-off. However, its performance becomes unstable as training progresses, eventually converging to the same final result as RL-zero. In contrast, BRIDGE demonstrates consistent improvement throughout training. Overall, nearly all hybrid baselines outperform RL-zero in terms of early-stage efficiency, highlighting the advantage of integrating supervised fine-tuning and reinforcement learning paradigms.
Method Average Performance Average
Epoch=1 Epoch=3 Epoch=6
@@
Table 4: Performance progression across training epochs for different methods.
Figure 3: Training dynamics of mean reward and response length for BRIDGE, Cold-start, and RL-zero on Qwen2.5-3B.
Training Dynamics Analysis. We analyze the dynamics of mean reward and response length during training for BRIDGE, Cold-start, and RL-Zero on Qwen2.5-3B. As shown in Figure 3, the three methods exhibit markedly different patterns. RL-Zero suffers from online RL’s sample inefficiency, showing slow growth in both response length and reward. Cold-start begins with extremely long responses due to SFT warm-up, causing slow initial training, followed by a sharp decline and gradual recovery. This "dip-then-rise" pattern indicates the model initially loses expert behavior acquired during SFT, then slowly explores new strategies—a mismatch that contributes to training inefficiency. Despite starting with higher rewards, Cold-start’s second-phase RL lacks proper guidance, resulting in convergence similar to RL-Zero. In contrast, BRIDGE benefits from continuous SFT guidance throughout training, enabling rapid reward growth that surpasses Cold-start and achieving superior convergence. These dynamics demonstrate that BRIDGE’s bilevel optimization enables more efficient policy learning through sustained and targeted expert guidance.
@@
Table 5: Cost-performance analysis on Qwen2.5-3B and Qwen3-8B-Base

### v2

@@ Abstract
Reinforcement learning (RL) has proven effective in incentivizing the reasoning abilities of large language models (LLMs), but suffers from severe efficiency challenges due to its trial-and-error nature. While the common practice employs supervised fine-tuning (SFT) as RL warmup, this decoupled two-stage approach suffers from catastrophic forgetting: second-stage RL gradually loses SFT-acquired behaviors and inefficiently explores new patterns. We introduce BRIDGE, a novel method to employ bilevel optimization to facilitate better cooperation between these training paradigms. By conditioning the SFT objective on the optimal RL policy, our approach enables SFT to meta-learn how to guide RL’s optimization process. During training, the lower-level performs RL updates while simultaneously receiving SFT supervision, while the upper-level explicitly maximizes the cooperative gain—the performance advantage of joint SFT-RL training over RL alone. Empirical evaluations across three LLMs and five reasoning benchmarks demonstrate that our method consistently outperforms baselines and achieves a better balance between effectiveness and efficiency. Specifically, BRIDGE achieves 44% faster training with a 13% performance gain on Qwen2.5-3B, and 14% faster training with a 10% improvement on Qwen3-8B.
Figure 1: Training dynamics of mean reward and response length on Qwen2.5-3B.
@@ 1 Introduction
The emergence of OpenAI’s o1 (OpenAI, ) and DeepSeek-R1 (DeepSeek-AI et al., 2025) marks a significant advance in LLM reasoning capabilities, particularly for challenging tasks such as mathematics (Cobbe et al., 2021; Hendrycks et al., 2021b) and programming (Chen et al., 2021; Codeforces, 2025). The key technique driving this progress is large-scale, rule-based RL. However, the inherently trial-and-error nature of RL renders the training process highly inefficient (RL-zero in Figure 1). An alternative approach is SFT on curated long chain-of-thought (CoT) datasets, which enables models to rapidly acquire reasoning patterns through imitation learning. While more efficient, SFT typically exhibits poorer performance and generalization than RL (Chu et al., 2025).
In practice, production-scale pipelines often use a two-stage cold-start paradigm, warming up with SFT before applying RL. This decoupled design introduces two issues: (a) catastrophic forgetting inherent to two-stage training, and (b) inefficient exploration in the second-stage RL due to the absence of guidance. As shown in Figure 1, the cold-start method exhibits a characteristic U-shaped trajectory in response length during the RL stage—first dipping and then rising—indicating that the model forgets previously learned expert patterns and explores new ones inefficiently. Meanwhile, the RL reward grows slowly due to the lack of guidance. As a result, the two stages fail to synergize, and both efficiency and performance remain unsatisfactory. This raises a natural question:
@@
To explore this possibility, we first propose a simple baseline that alternates between SFT and RL updates during training (Algorithm 1). Despite its simplicity, this approach improves both convergence efficiency and final performance (see Figure 2). However, such independent updates cannot guarantee improvements over RL alone, as not all SFT updates benefit RL optimization. Building on this insight, we develop BRIDGE, a cooperative learning framework based on bilevel optimization, where SFT serves as the upper-level problem and RL as the lower-level problem. By solving this nested structure—with the SFT objective explicitly conditioned on the RL solution—SFT provides targeted guidance that directly supports RL’s optimization process.
@@
To validate the effectiveness of our approach, we conduct experiments with three LLMs across five diverse benchmark datasets covering both standard and competition-level math reasoning tasks. Results demonstrate that BRIDGE consistently outperforms all baselines—including SFT, RL-zero, cold-start, and our alternating baseline—while requiring less wall-clock training time. These improvements confirm the benefits of tightly coupling SFT and RL through bilevel optimization rather than treating them as separate phases.
@@ 2.1 Fine-tuning Methods for Reasoning Models
We consider a language model parameterized by $\bm{\theta}$, which defines a conditional distribution $\pi(y|x;\bm{\theta})$ over output sequences $y$ given input sequences $x$. This work focuses on three widely used methodologies for fine-tuning $\bm{\theta}$ to enhance the model’s reasoning capabilities.
@@
Reinforcement learning with verifiable rewards. RLVR has gained increasing attention for its effectiveness in training advanced reasoning models such as DeepSeek-R1 (DeepSeek-AI et al., 2025). Given a dataset $\mathcal{D}_{\mathrm{RL}}:=\{(x,y)\}$ with verifiable outputs—such as mathematics competition problems—the objective of rule-based RL is formulated as:
@@
Here, $y$ denotes the ground-truth answer and $\hat{y}$ is the model’s predicted output. The equivalence relation $\hat{y}\equiv y$ is typically computed by a rule-based verifier. This objective is commonly solved using policy optimization methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (DeepSeek-AI et al., 2025).
@@ 2.3 A Simple Alternating Baseline
As shown in Figure 2, this alternating strategy converges faster than pure RL and achieves better final performance than both standalone SFT and two-stage cold-start training. While this integration yields empirical gains, the current formulation treats SFT and RL as independent update processes with no guarantee that alternating updates will consistently outperform RL method alone. This limitation prompts us to investigate: Can we develop a principled optimization framework where SFT and RL updates are inherently coordinated to guarantee improvements over pure RL?
@@ 3.1 BRIDGE: Cooperative Meta-Learning for SFT and RL
Given an SFT dataset $\mathcal{D}_{\mathrm{SFT}}$ and an RL dataset $\mathcal{D}_{\mathrm{RL}}$ (defined in Section 2.1), our objective is to integrate policy optimization (Eq. equation 2) with supervised learning (Eq. equation 1). We propose the following cooperative meta-learning formulation:
@@
For clarity, we express Equation equation 4 in simplified notation:
@@
The augmented model architecture, comprising base model parameters $\theta$ and LoRA parameters $w$, is essential for enabling cooperative learning. This separation allows the upper- and lower-level objectives to co-adapt during training, as illustrated in Figure 3. Without this architectural separation, our formulation (Equation equation 4) would collapse to a Model-Agnostic Meta-Learning (MAML)-style setup (Finn et al., 2017), where the lower-level solution reduces to a single gradient step used to update the upper-level SFT parameters. In this case, RL learning is disabled, and the cooperation between SFT and RL is lost.
@@ 3.2 Learning Algorithm
To solve the bilevel optimization problem in Eq. equation 5, we employ penalty-based methods (Shen & Chen, 2023; Shen et al., 2025) to avoid expensive second-order derivative computations. We first reformulate equation 5 as a single-level problem amenable to efficient first-order optimization.
@@
Algorithm 2 presents the learning procedure for BRIDGE. At each iteration, we sample SFT and RL mini-batches; update the base parameters $\theta$ with the joint objective; update the auxiliary parameters $\hat{\theta}$ via pure RL to track a baseline; and optimize the LoRA parameters $w$ to maximize cooperative gain—the improvement of the joint objective over pure RL. This process lets SFT meta-learn how to guide RL’s optimization, mitigating catastrophic forgetting and inefficient exploration.
@@ 3.3 Intuition Behind the Update Rules
Lower-level update: Gradient fusion. The update rule for $\theta$ in Eq. equation 8 performs a convex combination of SFT and RL gradients. As $\lambda$ increases from 0 to 1 during training, the algorithm smoothly transitions from imitation learning to reinforcement learning.
Upper-level update: Maximizing cooperative gain. The update for $w$ in Eq. equation 11 solves the bilevel problem by finding LoRA parameters $w$ that ensure the RL-optimized model $\theta^{*}(w)$ also excels on the supervised dataset $\mathcal{D}_{\mathrm{SFT}}$.
The update in Eq. equation 11 can be interpreted as performing gradient ascent on the following objective:
@@ 4.1 Settings
We use LIMR (Li et al., 2025) (1.3k problems; Qwen2.5-3B) and MATH (Hendrycks et al., 2021a) (8.5k; lama-3.2-3B-Instruct and Qwen3-8B) for RL training For the SFT dataset, we pair queries from LIMR and MATH with corresponding intermediate reasoning traces extracted from DeepMath-103k (He et al., 2025), which were distilled from the DeepSeek-R1 model. We evaluate on five mathematical reasoning benchmarks: MATH500 (Hendrycks et al., 2021a), Minerva Math (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), AIME 2024, and AMC 2023, and two out-of-distribution (OOD) benchmarks (LiveCodeBench v5  (Jain et al., 2024) and GPQA (Rein et al., 2024)) to assess generalization.
@@
To demonstrate the generality of our approach, we experiment with three LLMs: Qwen2.5-3B (Yang et al., 2024), Llama-3.2-3B-Instruct (Grattafiori et al., 2024), and Qwen3-8B-Base (Yang et al., 2025). All models use prompt formats consistent with SimpleRL (Zeng et al., 2025).
We use Verl  (Sheng et al., 2024) for RL training; full settings are in Appendix B.
@@ 4.2 Baselines
We compare BRIDGE against five baselines on the same base models:
@@
Base or instruction-tuned model with no reasoning-specific training.
@@
Trained only on curated reasoning traces by SFT (no RL).
@@
RL applied to the base model from scratch (no SFT warm-up).
@@
Two-stage SFT then RL, with fully decoupled objectives.
@@
Alternates SFT and RL updates without a cooperative objective.
@@ 4.3 Experimental Results
Table 1: Performance of BRIDGE compared to baselines on Qwen2.5-3B across five math benchmarks Average relative performance gains (%) over Cold-start are highlighted in blue.
@@
Naive Alter. 65.2 25.3 27.1 6.7 42.5 33.4 (+3.7%)
BRIDGE 66.2 23.9 28.9 13.3 47.5 36.0 (+11.8%)
Table 2: Performance on Llama3.2-3B-Instruct.
@@
Main results across three LLMs. Across five math benchmarks and three LLMs (Tables 1, 2, 3), BRIDGE attains the highest average accuracy. Cold-start behaves inconsistently across backbones: on Llama-3.2-3B-Instruct it is lower than RL-zero on average, whereas on Qwen3-8B-Base it exceeds RL-zero; this suggests that the two stage SFT then RL pipeline can constrain subsequent exploration and is not reliably optimal. Naive Alternating exceeds RL-zero and Cold-start on Qwen2.5-3B and Qwen3-8B-Base, indicating that training SFT and RL at the same time helps, although the gains are limited. In contrast, BRIDGE consistently outperforms all methods on all three backbones, with average improvements over Cold-start of 11.8% on Qwen2.5-3B, 30.9% on Llama-3.2-3B-Instruct, and 9.7% on Qwen3-8B-Base. Unlike naive alternating, BRIDGE treats SFT and RL as a cooperative objective and adapts the supervision so that it benefits RL objective; this coupling yields robust and transferable gains across instruction tuned and base models and remains effective at larger scales.
Generalization to More Challenging Math Tasks Baseline methods tend to yield larger improvements on relatively easier benchmarks but generalize poorly to more complex reasoning tasks. For example, the Cold-start underperforms RL-zero on Minerva Math, OlympiadBench, and AMC23, indicating that Cold-start can restrict exploration and transfer poorly to harder problems. While the Naive Alternative partially mitigates this issue, its gains remain limited. In contrast, BRIDGE achieves consistent and substantial improvements on the more challenging benchmarks. These results underscore BRIDGE’s superior generalizability in handling complex mathematical reasoning.
Table 3: Performance on Qwen3-8B-Base.
@@
Generalization to Out-of-Domain tasks. To test whether BRIDGE’s benefits transfer beyond math into out-of-domain reasoning. We evaluate on two non-math OOD benchmarks, LiveCodeBench and GPQA based on Qwne3-8B, as seen in Table 4.3, BRIDGE attains the best performance, demonstrating that joint training yields consistent cross-domain gains.
Table 4: OOD evaluation on Qwen3-8B.
Method LiveCodeBench GPQA Avg.
Base 32.95 32.32 32.64
Instruct 33.07 35.80 34.44
RL-zero 32.50 38.38 35.44
Cold-start 23.86 25.76 24.81
BRIDGE 34.55 42.93 38.74
Table 5: Average performance across epochs.
Method Epoch=1 Epoch=3 Epoch=6 Avg.
@@
Training Dynamics Analysis. We analyze the dynamics of mean reward and response length during training for BRIDGE, Cold-start, and RL-Zero on Qwen2.5-3B. As shown in Figure 1, the three methods exhibit markedly different patterns. RL-Zero suffers from online RL’s sample inefficiency, showing slow growth in both response length and reward. Cold-start begins with extremely long responses due to SFT warm-up, causing slow training (see in Table 6), followed by a sharp decline and gradual recovery. This "dip-then-rise" pattern indicates the model initially loses expert behavior acquired during SFT, then slowly explores new strategies—a mismatch that contributes to training inefficiency. Despite starting with higher rewards, Cold-start’s second-phase RL lacks proper guidance, resulting in final rewards similar to RL-Zero. In contrast, BRIDGE benefits from continuous SFT guidance throughout training, enabling rapid reward growth that surpasses Cold-start and achieving superior convergence. These dynamics demonstrate that BRIDGE’s bilevel optimization enables more efficient policy learning through sustained and targeted expert guidance.
Performance on downstream tasks across training epochs shows the same pattern (Table 4.3): Cold-start learns quickly at the beginning but dips mid-training and finishes no better than RL-zero, indicating constrained exploration. BRIDGE starts strong and improves steadily to the best final performance. Naive Alternating narrows the gap mid-training yet remains below BRIDGE. RL-zero shows the poorest early-stage efficiency. (see Appendix C for details.)
@@
Table 6: Cost-performance analysis on Qwen2.5-3B and Qwen3-8B-Base
@@
LoRA ablation. We prove BRIDGE is insensitive to LoRA hyperparameters (Appendix D).